If 2 workers can assemble a car in 8 hours and third workers can assemble the car in 12 hours, then how long will take the 3 workers to assemble the car?
Answer by ptaylor(2198)


Let x=amount of time needed to assemble the car if all three workers are working together
Then all three workers work at the rate of 1/x car assembled per hour
Two workers work at the rate of 1/8 car assembled per hour
Third worker works at the rate of 1/12 car assembled per hour
Then the three workers work at the rate of 1/8 + 1/12 car assembled per hour, so:
1/8 +1/12 =1/x multiply each term by 24x
3x+2x=24
5x=24
x=4.8 hours --time it takes to assemble the car with all three working together
CK
In 4.8 hours the two workers assemble (1/8)*4.8=0.6 of the car
In 4.8 hours the third worker can assemble (1/12)*4.8=0.4 of the car
0.6+0.4=1 (1 car, that is)
Hope this helps---ptaylor